How do I graph the following using slope?
1. 4x + y = 6
2. 2y -2 = 6x
3. 3y - 6 = 6x
4. 2x + 6y = 12

Answer by stanbon(75887)

How do I graph the following using slope?
----------------------
In each case you start with the y-intercept, then you apply the slope.
----------------
1. 4x + y = 6
y =-4x + 6
1st: Mark the intercept at (0,6)
2nd: The slope is -4/1
Procedure: move your pencil 1 to the right and then down 4 to
the point (1,2). Draw a line thru the two points.
----------------------
2. 2y -2 = 6x
y = 3x + 1
1st: Mark (0,1)
2nd: The slope is 3/1
Move your pencil 1 to the right and up 3 to the point (1,4)
Draw a line thru the two points.
------------------------
3. 3y - 6 = 6x
y = 2x + 2
1st: mark (0,2)
2nd: the slope is 2/1
move your pencil 1 to the right and then up 2 to the point (1,4)
Draw the line.---------------------------
4. 2x + 6y = 12
y = (-1/3)x + 2
slope is (-1/3) which can be written as (-1/3)/1
intercept = (0,2)
mark (0,2)
move your pencil 1 to the right and down (1/3) to the point (1,1 2/3)
or (1,5/3)
Draw the line
=====================
Cheers,
Stan H.